# Continual Release Moment Estimation with Differential Privacy

### v1

@@ C.1 Bounds on the expected approximation error for PP with non-trivial factorization
$S_{2}$ $=\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}=\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}(A_{2})_{k,t}(C_{1}^{-1}Z_{1})^{2}_{t,j}\right)^{2}$
@@
$\mathbb{E}(C_{1}^{-1}Z_{1})^{2}_{t_{1},j}(C_{1}^{-1}Z_{1})^{2}_{t_{2},j}$ $=\mathbb{E}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}(Z_{1})_{r,j}^{4}+\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}$
@@
$=3\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}+\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}$
@@
$\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+2Q_{t_{1},t_{2}}^{2})$
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}_{Z}\|\widehat{D}_{\text{PP}}-D\|^{2}$ $=2d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))+4\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{\top})$
@@
$[\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))]_{k,j}$ $=\mathbb{E}\sum\limits_{t=1}^{n}(A_{2})_{k,t}\left(\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,j}\right)^{2}$
@@
$\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}$
@@
$\mathbb{E}\langle A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1})),\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\rangle=\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}.$
@@
Appendix D Algorithms
Algorithm 2 Differentially Private JME Adam
  $m_{0}\leftarrow 0$   // first moment initialization.
  $v_{0}\leftarrow 0$   // second moment initialization.
  for $i=1,2,\dots,n$ do
     $S_{i}\leftarrow\{d_{1},\dots,d_{m}\}\subseteq D$ select a data batch
     $x_{i}\leftarrow\sum_{j=1}^{m}\min(1,\zeta/||g_{j}||)g_{j}$
     $\widehat{x}_{i}\leftarrow x_{i}+\zeta[C_{\beta_{1}}^{-1}Z_{1}]_{[i,\cdot]}$
     $m_{i}\leftarrow m_{i-1}\beta_{1}+(1-\beta_{1})\widehat{x}_{i}$
     $v_{i}\leftarrow v_{i-1}\beta_{2}+(1-\beta_{2})\widehat{x^{2}}_{i}$
     $\widehat{m}_{i}=m_{i}/(1-\beta_{1}^{i})$   // bias-correction
     $\widehat{v}_{i}=v_{i}/(1-\beta_{2}^{i})$   // bias-correction
  end for
  $\Theta=(\theta_{1},\dots,\theta_{n})$
Algorithm 3 $\lambda$-JME
0:  input stream vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
0:  workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
0:  noise shaping matrices $C_{1},C_{2}$ (lower triangular, invertible, decreasing column norms)  (default: $\text{I}_{n\times n}$)
0:  privacy parameters $(\epsilon,\delta)$
  for $t=1,2,\dots,n$ do
     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$
[…]
  end for
0:  input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
0:  workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
0:  noise shaping matrices $C_{1},C_{2}$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
0:  privacy parameters $(\epsilon,\delta)$
0:  privacy trade-off $\alpha\in(0,1)$
  $(\epsilon_{1},\delta_{1})\leftarrow(\alpha\epsilon,\alpha\delta)$ // 1st moment private budget
  $(\epsilon_{2},\delta_{2})\leftarrow((1-\alpha)\epsilon,(1-\alpha)\delta)$ // 2nd moment private budget
  $\sigma_{1}\leftarrow$ noise strength for $(\epsilon_{1},\delta_{1})$-dp Gaussian mechanism
  $s_{1}\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // sensitivity of 1st moment
  $\sigma_{2}\leftarrow$ noise strength for $(\epsilon_{2},\delta_{2})$-dp Gaussian mechanism
  for $t=1,2,\dots,n$ do
     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$
[…]
  end for
Algorithm 5 $\tau$-CS (CS with 2nd moment rescaling parameter $\tau>0$)
0:  input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with
0:  workload matrices $A=(a^{t}_{k})\in\mathbb{R}^{n\times n}$
0:  input dimension $d$, bound on input norm $\zeta>0$
0:  privacy parameters $(\epsilon,\delta)$
0:  (optional) noise shaping matrix $C$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
  for $t=1,2,\dots,n$ do
[…]
  end for
Algorithm 6 PP
0:  input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
0:  workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
0:  noise shaping matrix $C_{1}$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
0:  privacy parameters $(\epsilon,\delta)$
  $s\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // sensitivity of 1st moment
  for $t=1,2,\dots,n$ do
     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z]_{[t,\cdot]}$
[…]
  end for

### v2

@@ D.1 Bounds on the expected approximation error for PP with non-trivial factorization
$S_{2}$ $=\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}=\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}(A_{2})_{k,t}(C_{1}^{-1}Z_{1})^{2}_{t,j}\right)^{2}$ (131)
@@
$\mathbb{E}(C_{1}^{-1}Z_{1})^{2}_{t_{1},j}(C_{1}^{-1}Z_{1})^{2}_{t_{2},j}$ $=\mathbb{E}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}(Z_{1})_{r,j}^{4}$ (132)
$\quad+\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}$
@@
$=3\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}$
$\quad+\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}$
@@
$\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+2Q_{t_{1},t_{2}}^{2})$ (133)
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}_{Z}\|\widehat{D}_{\text{PP}}-D\|^{2}$ $=2d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))$ (134)
$\quad+4\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{\top})$
@@
k,j $=\mathbb{E}\sum\limits_{t=1}^{n}(A_{2})_{k,t}\left(\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,j}\right)^{2}$ (135)
@@
$\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}$ (136)
@@
$\mathbb{E}\langle A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1})),\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\rangle=\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}.$ (137)